OpenStudy (jnavao):

how to solve 3 variable system ? Help please will give a medal!!

OpenStudy (mayankdevnani):

For simplest way to solve 3 variable system is use DETERMINANT METHOD .

OpenStudy (mayankdevnani):

Welcome to OPEN STUDY btw @jnavao

OpenStudy (michaelgrabiski):

You need to think of which variable as a plane in 3 dimensions so like x, y, and z so then you take the first two equitation and add them and you take the last two which means your reusing the 2nd equation. here's an example x-y+2x=5 3x+y-z=-1 5x-1y+2z=19 so take the first to and you get 4x-z=4 and the 2nd two and you get 8x+z-14 then take those two and you get 12x/12 =24/12 which is then x =2 then substitute the x back in to find the other variables watch this - https://www.khanacademy.org/math/algebra2/advanced-equations-and-inequalities/systems-with-three-variables/v/systems-of-three-variables

sam (.sam.):

You are currently in Openstudy Feedback section. Please post in the right section: http://openstudy.com/study#/groups/Mathematics